My question is:
Find the length of the diagonal of a rectangle whose length is 8m and width is 5m.
thanks

Answer by arunpaul(104)

length = 8 m
width = 5m
so the length of diagonal = l
so as perPythagorean-theorem
l^2 = 8^2+5^2=64+25=89
so l = SQRT 89= 9.433m app (Ans)